OpenStudy (anonymous):

Maria has the following scores on exams in her social studies class: 87, 77, 95, 59, 92, and 59. a. Find the mean, median, and mode of the scores. b. Which measure (mean, median, or mode) should Maria's social studies teacher use to convince Maria that she needs to study more regularly for her exams? \

OpenStudy (anonymous):

A) First of all, line up the scores from least to greatest: 59, 59, 77 87, 92,95 For the median, find the number in the center. Make sure your numbers are in order! MEDIAN = 81 For the mean, add up all the numbers and divide that by the amount of numbers. In this case, add up 87, 77, 95, 59, 92, and 59. Then, divide that number, 469, by 6. 469/6 = 78.1 MEAN = 78.1 The mode is the number that occurs the most, in this case 59, which appears twice. All of the other numbers only appear once. MODE = 59 B) Maria's social studies teacher should use the mode, because its the lowest number. This will convince Maria to study more regularly for her exams hoope this helps :)

OpenStudy (anonymous):

@puper15

OpenStudy (kropot72):

Shouldn't the median be the arithmetic mean of 77 and 87 = 164/2 = 82

OpenStudy (anonymous):

median is 77 mean is 75.4 mode 59